OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 69.

OpenStudy (anonymous):

x(x+1) = 69

OpenStudy (anonymous):

x + x + 1 = 69

OpenStudy (anonymous):

x+(x+1)=69